OpenStudy (anonymous):

what is the integral of f(x) when f(x)=sqrt((3-x)/(x+1)) please show all steps thank you!!

OpenStudy (chris911):

lol wassup its early O,O

OpenStudy (anonymous):

what is the integral of f(x) when f(x)=sqrt((3-x)/(x+1)) please show all steps thank you!!

OpenStudy (anonymous):

@AccessDenied @mathavraj @Hero @SolomonZelman

OpenStudy (anonymous):

@mathavraj

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

I am not good at integrating (yet) -:(

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

someone help

OpenStudy (anonymous):

i think u have to do it by substitution method

OpenStudy (solomonzelman):

idk, I would guess you multiply top and bottom times √(x+1) to rationalize the denominator. Will see what accessdenied says....

OpenStudy (accessdenied):

I am imagining an interesting u-substitution. If you chose u = sqrt(x + 1), you could use some algebra to put this in terms of sqrt(3 - x) equaling a function of u. And taking the du part will give you what is in the denominator as well. Can you see what I am going for here? And if so, what do you get by trying that route?

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

seond one

OpenStudy (anonymous):

ty one more question

Parth (parthkohli):

Weren't you in 8th?

OpenStudy (anonymous):

@ParthKohli THAT WAS MY YOUNGER BROTHER QUESTION HE USES MY ACCOUNT

hartnn (hartnn):

i am not sure how u= u=x+1 would work... i was thinking more on the lines of \(\Large \dfrac{3-x}{\sqrt{(3-x)(x+1)}}\) then separate the denominator

hartnn (hartnn):

no, don't separate the denominator....

hartnn (hartnn):

arrange the numerator in terms of the DERIVATIVE of denominator and then separate it out

OpenStudy (anonymous):

THX now answwer my younger brothers question

hartnn (hartnn):

whats your younger bro's question ?

OpenStudy (anonymous):

the present age of mr jackson son is in the ratio 17:9 if the ratio 9 years ago was 7:3 find there present age

OpenStudy (anonymous):

its a ratio sum

hartnn (hartnn):

let their present age be x and y ratio = 17/9 so x/y = 17/9

hartnn (hartnn):

after 9 years... their ages will be x+9 y+9

hartnn (hartnn):

ratio then will be (x+9)/(y+9) =7/3 now you have 2 equations 2 unknowns

OpenStudy (anonymous):

ty @hartnn

OpenStudy (marissalovescats):

How is it set up? Like this?